OpenStudy (anonymous):

What is the value of the coefficient of "a" when the quadratic equation y = (2x − 1)(4x − 3) is written in standard form? 6 8 −10 −6

OpenStudy (anonymous):

(2x-1)(4x-3)=8x^2-10x+3(standard form) so a=8 in ax^2 + bx + c = 0